Sol Summary – November 20th

Sol: 7
Summary Title: Biologist down
Author’s name: Jas Purewal
Mission Status: Nominal
Sol Activity Summary:
Read about lichen and did research on how to identify lichen. Planted 18 plants. Commander, Communications officer and Engineer did a walking EVA at 2:45pm to Hab Ridge Rd N. Searching for hydrous minerals. Conducted an emergency exercise that simulated a crew member fainting. The crew performed admirably.
Look Ahead Plan: Morning crisis communication education and VR first aid training and scenario 3. 1:30pm to 4:30pm EVA to Sea of Shells via the Overlook on Brahe Hwy 1572.
Anomalies in work: None
Weather: Rather nice. Sunny, clear. No rain.
Crew Physical Status: Good
EVA: EVA 8 – Commander, Communications officer and Engineer did a walking EVA at 2:45pm to Hab Ridge Rd N. Searching for hydrous minerals. Found many sea shells, they were not collected but rock samples were collected.
Reports to be filed:
Commander’s report
EVA report
EVA request
Operations report
Sol Summary report
Journalist report

Support Requested: none

Operations Report – November 20th

SOL: 7
Name of person filing report: Jas Purewal
Non-nominal systems: Drain under kitchen sink leaked again, but we have a temporary fix.
Notes on non-nominal systems: NA
Spirit rover used: No
Hours: (before EVA) NA
Beginning charge: (Before EVA) NA
Ending charge: (On return from EVA, before recharging) NA
Currently charging: Yes
Opportunity rover used: No
Hours: NA
Beginning charge: NA
Ending charge: NA
Currently charging: NA
Curiosity rover used: No
Hours: NA
Beginning charge: NA
Ending charge: NA
Currently charging: NA
Perseverance rover used: No
Hours: NA
Beginning charge: NA
Ending charge: NA
Currently charging: NA
General notes on rovers: NA
Summary of Hab operations: put notes here
WATER USE: 50 gallons
Water (static tank): 250 gallons
Static tank pipe heater (on or off): on
Static tank heater (On or off) off
Toilet tank emptied: No
Summary of internet: Good
Summary of suits and radios: optimal
Summary of GreenHab operations: None
WATER USE: 6 gallons
Heater: On
Supplemental light: Off
Harvest: None
Summary of ScienceDome operations: None
Dual split: (Heat or AC, On or Off) Heat, Off
Summary of RAM operations: (Tools used, work done) Plumbing adhesive and tape currently in main hab. Took needle nose pliers from the RAM, currently in the main hab.
Summary of any observatory issues: NA
Summary of health and safety issues: none
Questions, concerns and requests to Mission Support: Still waiting for pillows. Whilst we are conserving water, at our current rate of use we predict that we will be down to our last 50 gallons by Sol 11.

Journalist Report – November 20th

Sol: 7
Title: Martian Sunday
Author: Izabela Shopova, Crew Journalist
We made it through our first week on Mars. And celebrated Sunday with a very relaxed plan for the day. Well, relaxed by the crazy standards of Mars. Our crew engineer Judi repaired the leaking sink in the kitchen with the help of the XO, Caitie, the crew biologist watered the plants and planted new seeds in the GreenHab, we all did house chores and wrote in our journals.

In the afternoon we did another EVA – this time without our faithful companions – the rovers. Instead of riding them on the road, we took off on foot and explored the area around the Hab, looking for hydrous minerals in the dried river beds and on top of the ridge. We found so many seashells, a whole martian sea of sea shells, crunching under our feet, piling up on the side of the hill. The sunset light made the canyons below look like a multi-layered cinnamon and strawberry cake.

Back in the safety of the Hab we were uploading photos after the EVA , when crew biologist Caitie went to the GreenHub to water the plants, but when we called her on the radio she didn’t respond. Jennie and I went to check on her and found our crew biologist lying on the ground. We immediately initiated an emergency response by calling Hab on the radio and assessing the unconscious crew member. Long story – short: Catie was assessed and brought to safety inside the Hab when the rest of the crew finally revealed that it was all just an emergency response simulation that they had planned in secrecy. They all played their roles with gusto and conviction, so it was a great relief to find out that Caitie was well. HSO provided us with glowing feedback for our appropriate response and we had a quick overview of the first aid protocol before dinner.

Sunday also marked the end of my time with Paro. At dinner we took our last photo together. From this evening Jennie will be spending time with him and I won’t be allowed to cuddle him anymore. Next week will be a tough one.

GreenHab Report – November 20th

Sol 7
GreenHab Officer: Caitlyn Hubric
Environmental control: (heater, fan, door open?) heater on, door closed, 20% humidity
Average temperatures: 43.3ºC/13.5ºC
Hours of supplemental light: N/A
Daily water usage for crops: 22.6 L ≈ 5.97 gal ≈ 6 gal
Daily water usage for research and/or other purposes: 0 gal
Water in Blue Tank ~265 gallons
Time(s) of watering for crops: twice, morning and night, 9am-5pm
Changes to crops:
-Today I replanted multiple crops that haven’t grown since I started watering the greenhouse. I replanted the mint in the single planter on the wood pallet, I replanted 5 beans (bean bush contender) in the single pots on the shelves (middle row), I planted spinach by the spinach sign in the blue bin, I replanted some radish in the spots that I harvested the other day, and I planted 2 peas (super snappy peas) in the single pots on the top layer of the shelves. I only planted things that weren’t growing and followed the label they already had, I didn’t bring in any new plants.
-The beans currently growing look very happy as do the sugar snap peas.
-I hope that the plants I replanted will grow successfully so the future crews can enjoy a variety of fresh plants.
Harvest: (include which crop and mass in grams) N/A
Support/supplies needed: N/A

EVA Report – November 20th

EVA # 8
Author: Izabela Shopova
Purpose of EVA: Searching for hydrous minerals — clays, sulfates, silica — and/or carbonate salts. Obtaining additional rocks for engineering demonstration (for use with mechanical anchor prototype). Exploring area nearby habitat under alternate lighting.
Start time: 2:45 PM
End time: 4:00 PM
The crew went to explore the area around the Hab and the top of Hab Ridge to search for hydrous minerals. We followed a dried river bed for the first part of the expedition, then turned west and climbed up the side of the ridge. We were able to identify and collect all 3 of the minerals we were searching for. And in addition, we came across a big deposit of seashells at the top of the ridge. After taking photos of the locations of our discoveries, we turned around and walked back to the Hab, taking a longer route to the east. We ended the EVA with a walk around the base and a visual check, confirming that the tunnels and external walls don’t show any new damage, before entering the airlock and completing the EVA.
Destination: Hab Ridge Road
Coordinates (use UTM WGS 84): 518000, 4251000 to 518000, 4250000
Participants: CDR, COMM, ENGR
Road(s) and routes per MDRS Map:
Walking north from Hab via Sagan Street (1103) to Hab Ridge Road North (518000, 4251000). Then south along Hab Ridge Road (518000, 4250000). Returning via the same path

Mode of travel: Walking

Commander Report – November 20th

Sol: 7
Summary Title: “Caitie’s Down!”
Author’s name: Dr. Jenni Hesterman, Commander
Mission Status: Nominal
Commander Report:

After an evening of card games, the crew decided to sleep in and enjoy a slow start to their Sunday. After brunch, three crew members set out on foot for Hab Ridge Road to search for water on Mars and located hydrous minerals in the form of clays, sulfates, silica and carbonate salts. At the top of the path, they were surprised to find mounds of seashells, likely left behind when the Western Inland Sea retreated from Utah during the late Cretacious period, 100-60 million years ago. Upon return, the crew accomplished a visual inspection of the Hab and found no damage or anomalies.

While the crew was cooking dinner, Crew Biologist Caitie went to the GreenHab to accomplish the evening watering. When she did not check in as expected, the Commander the Crew Journalist went to her location for a health and safety check. We found her lying on the ground outside the GreenHab. She was responsive and said she may have passed out and wasn’t sure if she hit her head. We covered her with our sweaters, and called the other crew members to inform them of the situation, asking for the Health and Safety Officer (HSO) to respond ASAP. The HSO quickly arrived and assessed the patient, determining she was stable and recommending relocation inside the Hab. Caitie was able to stand and walk under her own power, with support from the crew.

After entering the Hab, we were informed by the XO, HSO and Caitie that this was an emergency response simulation they designed to test our crew’s crisis response protocols. We are at the halfway point of the mission and settled into our routines, so the exercise was a good reminder to fight complacency. When teams get comfortable with a high level of performance, overconfidence can lead to massive failures. We are thankful for this opportunity to hone our skills and competencies.

Copyright © The Mars Society. All rights reserved. | Main Site